# Searching from Area to Point: A Semantic Guided Framework with Geometric Consistency for Accurate Feature Matching

### v5

@@ IV-C Framework Implementation
$\mathcal{P}_{g}^{C}=GMC_{PM}(\{\mathcal{A}_{i,\pi(i)},\mathcal{P}_{i}\}_{i}^{T},I_{0},I_{1},T_{SP})$ (14)
Finally, the point matches including inside-area matches $\mathcal{P}_{i}\}_{i}^{T}$ and global matches $\mathcal{P}_{g}^{C}$ are achieved as the output of our SGAM method, which possess both high matching accuracy and uniform spatial distribution. It is noteworthy the $PM$ we utilized can be any detector-free method without re-training or fine-tuning. Therefore, our SGAM is a plug-and-play framework to improve the feature matching precision.
V Results
TABLE II: Value results (%) of MMA. We report MMA with three thresholds under various matching difficulties. Our SGAM is applied on four baselines. To show the impact of semantic accuracy to our method, we take three different semantic inputs: SGAM using ground truth (GT), SGAM using SEEM-L and SGAM using SEEM-T. The improvement achieved by SGAM is also reported in percentage, which is impressive to show the effectiveness of our methed.
MMA ScanNet: FD@5 ScanNet: FD@10 MatterPort3D
@@ V Results
SP [5]+NN 26.56 59.87 75.75 13.76 40.47 59.49 3.46 17.41 32.43
SP [5]+SG [30] 37.54 63.06 76.15 24.40 43.32 57.57 13.77 21.66 29.95
ASLFeat [9]+NN 28.90 58.69 75.52 15.66 40.98 60.41 6.31 20.59 35.79
ALike [8]+NN 31.99 60.98 75.03 16.41 40.98 59.49 7.25 20.94 34.55
ASpan [7] 32.99 66.91 85.03 25.35 49.83 70.79 7.17 21.10 37.25
@@
SEEM-L[16]+SGAM_ASpan 36.48+10.58% 70.70+5.66% 87.58+2.99% 27.15+7.09% 52.85+6.07% 73.76+4.19% 7.61+6.11% 23.98+13.66% 39.16+5.12%
SEEM-T[16]+SGAM_ASpan 35.54+7.73% 69.44+3.79% 86.64+1.89% 26.81+5.78% 52.11+4.59% 72.84+2.89% 7.40+3.17% 22.51+6.69% 38.41+3.11%
[…]
@@
SEEM-L[16]+SGAM_QuadT 37.02+12.91% 73.63+4.59% 89.30+1.12% 25.17+11.03% 60.55+6.38% 81.08+3.35% 7.91+6.41% 25.95+8.26% 43.05+3.17%
SEEM-T[16]+SGAM_QuadT 36.35+10.85% 72.54+3.04% 88.40+0.10% 24.30+7.15% 59.38+4.32% 80.46+2.55% 7.86+5.63% 24.50+2.23% 42.88+2.76%
[…]
@@
SEEM-L[16]+SGAM_LoFTR 33.83+10.95% 70.05+7.23% 87.33+4.58% 18.85+5.60% 48.78+4.27% 68.90+1.47% 12.27+29.16% 27.20+23.22% 40.25+11.57%
SEEM-T[16]+SGAM_LoFTR 33.17+10.55% 69.52+6.40% 86.71+3.84% 18.21+2.03% 47.45+1.42% 67.98+0.12% 11.47+20.77% 25.10+13.69% 38.45+6.59%
[…]
@@
SEEM-L[16]+SGAM_COTR 36.54+11.00% 66.48+4.78% 81.04+2.96% 18.16+10.01% 44.54+5.15% 63.29+3.76% 11.73+10.40% 31.97+8.88% 48.70+5.71%
SEEM-T[16]+SGAM_COTR 36.05+9.52% 65.89+3.85% 80.48+2.24% 17.60+6.60% 43.71+3.18% 62.50+2.47% 11.11+4.57% 31.30+6.58% 47.49+3.08%
V-A Dataset
@@
To demonstrate the superiority of the A2PM framework and SGAM method, we evaluate our methods on two different indoor datasets, ScanNet [43] and MatterPort3D [44]. Additionally, we investigated the robustness of our method in diverse semantic scenes by conducting experiments on the outdoor KITTI360 dataset [45] and YFCC100M [46] dataset. First three datasets all offer ground truth semantic labels, which can be directly used as the input of our method. Moreover, we also evaluated the robustness of our method against semantic prior precision by utilizing recent semantic segmentation method, SEEM [16], to perform semantic segmentation. The ScanNet contains numerous sequence images, and we selected image pairs with varying levels of difficulty based on the frame difference from its $scene\_0000$ to $scene\_0299$ to evaluate our method. We also compare with other SOTA methods on the standard ScanNet-1500 benchmark [30], which offers no semantic labels and we achieve the semantic prior by SEEM. Due to the data collection settings of MatterPort3D, image pairs with overlap in this dataset result in wide baseline and present challenging matching conditions. These conditions allow us to effectively showcase the performance of our method under difficult matching conditions. The KITTI360 dataset allows for the evaluation of driving scenes, which is a widely-used dataset for SLAM. The YFCC100M dataset contains internet images of architectural scenarios, which is also widely-used for feature matching.
V-B Implementation Details
1) Parameter Setting. In the SAM, two semantic object areas with center distance (Sec. IV-A1) less than $100$ pixels are fused. The parameter is set for sparse detection results. The multiscale ratios are $[0.8,1.2,1.4]$ for the scale invariance enhancement of two area descriptors, which aim to achieve more semantic specificity. The thresholds in $SOA$ and $SIA$ matching are set as $T_{\mathsf{H}}=0.5$, $T_{\mathsf{L}}=0.75$ and $T_{da}=0.2$, which are smaller for more restricted matching. In SIA detection, the top layer reduce ratio in semantic pyramid is $r\!=\!8$, which is a trade-off between detection efficient and accuracy. We empirically set the $\phi$ for $T_{GR}$ as $0.5$ for ScanNet and $1.0$ for other datasets, the ablation study for which can be seen in Sec. V-G. The $T_{SP}$ is set as $0.6$ for ScanNet and $0.3$ for other datasets, based on ablation study in Sec. V-H.
2) Area Size. The area size is the input size for the point matchers. In practice, the default area size is set as $256\times 256$ for ScanNet, $480\times 480$ for Matterport3D, YFCC, KITTI and $640\times 640$ for ScanNet1500. For the matched semantic object areas (SOAs), the sizes are first expanded from the bounding boxes of objects to match the width-height ratio of the default size. Then the SOAs are cropped from the original images and resized to the default size before being entered into the point matcher. After matching, the correspondences outside the object bounding boxes are filtered out. For semantic intersection areas (SIAs), their size is related to the detection window, which is first set with the default size. To further mitigate the scale issue between images, we first match SOAs, which are robust to scale variation as the actual sizes of objects are fixed. Then, we adjust the detection windows of SIA in two images using the average size variation of SOA bounding boxes to obtain areas with consistent scales.
3) Semantic Noise Filtering. Our method takes semantic segmentation images as input, obtained from SOTA semantic segmentation methods or ground truth labels. However, even if provided with manual labels, these inputs can still contain semantic labeling errors. Thus, the semantic noise filtering needs to be performed in SAM. Specifically, in semantic object area detection, objects smaller than ${1}/{100}$ image size is ignored and the semantic surrounding descriptor neglects semantics with fewer than $20$ continuous pixels at the boundary. In semantic intersection area detection, semantics smaller than ${1}/{64}$ window size are filtered. The semantic labels with size less than ${1}/{64}$ area size are filtered out in the construction of semantic proportion descriptor.
@@ V-B Implementation Details
TABLE III: Relative pose estimation Results (%). The AUC of pose error on ScanNet (FD@5/10) and MatterPort3D with different thresholds are reported. Our SGAM is applied on four baselines. To show the impact of semantic accuracy to our method, we takes three different semantic inputs: SGAM using ground truth (GT), SGAM using SEEM-L and SGAM using SEEM-T. The improvement achieved by SGAM is also reported in percentage.
Pose estimation AUC ScanNet: FD@5 ScanNet: FD@10 MatterPort3D
@@
SP [5]+NN 63.07 70.83 85.41 53.19 64.46 73.99 13.82 24.09 35.26
SP [5] +SG[30] 67.46 76.46 86.61 53.11 64.47 73.58 16.39 29.54 37.61
ASLFeat [9]+NN 68.08 75.64 85.66 56.89 68.66 78.33 12.64 25.91 39.26
ALike [8]+NN 69.67 78.29 83.46 50.29 61.27 70.31 15.14 20.03 23.79
ASpan [7] 70.73 77.41 80.19 58.51 70.42 79.84 18.35 27.81 43.98
@@
SEEM-L [16]+SGAM_ASpan 72.32+2.24% 80.59+4.10% 85.20+6.24% 59.17+1.13% 73.02+3.69% 83.40+4.46% 19.74+7.56% 29.38+5.66% 45.52+3.51%
SEEM-T [16]+SGAM_ASpan 71.84+1.56% 79.94+3.26% 83.32+3.90% 58.87+0.62% 72.02+2.27% 82.54+3.38% 18.64+1.56% 28.87+3.81% 44.34+0.83%
[…]
@@
SEEM-L [16]+SGAM_QuadT 71.92+3.51% 75.60+1.81% 81.43+2.56% 61.78+4.23% 71.96+3.14% 76.52+2.08% 17.83+7.82% 27.25+1.00% 41.56+4.00%
SEEM-T [16]+SGAM_QuadT 71.47+2.86% 75.06+1.09% 80.66+1.59% 60.97+2.87% 71.77+2.86% 75.80+1.13% 16.84+1.88% 27.10+0.44% 40.22+0.65%
[…]
@@
SEEM-L [16]+SGAM_LoFTR 70.25+3.79% 79.23+6.65% 83.93+6.99% 60.37+2.83% 71.02+1.74% 80.91+2.42% 18.04+0.31% 30.86+11.04% 42.13+10.32%
SEEM-T [16]+SGAM_LoFTR 69.53+2.72% 78.03+5.03% 82.32+4.93% 59.01+0.50% 70.58+1.10% 79.78+1.01% 17.47-2.85% 30.00+7.94% 40.63+6.39%
[…]
@@
SEEM-L [16]+SGAM_COTR 69.67+4.12% 78.15+5.46% 83.84+6.83% 52.80+1.70% 66.16+4.42% 78.78+8.59% 18.12+1.82% 26.99+7.62% 36.69+7.67%
SEEM-T [16]+SGAM_COTR 69.40+3.72% 77.98+5.32% 83.07+5.85% 52.07+0.28% 65.91+4.03% 78.43+8.10% 17.93+0.71% 25.79+2.84% 35.79+5.01%
V-C Point Matching
@@
We conducted point matching experiments on ScanNet and MatterPort3D to evaluate matching precision. For ScanNet, we construct two matching difficulties with image pairs under various Frame Differences (FD@5/10), each including 1500 image pairs. For the more challenging matching condition, 500 image pairs are sampled from 5 scenes in MatterPort3D.
@@
1) Compared methods. We compare the proposed A2PM framework with conventional matching frameworks including detector-based matching [30, 5, 9, 8] and detector-free matching [7, 11, 10, 37]. For A2PM framework, we combine SGAM with SOTA Transformer-based detector-free methods: SGAM_ASpan [7], SGAM_QuadT [11], SGAM_LoFTR [10] and SGAM_COTR [37]. To evaluate the robustness against semantic precision of our method, we provide our method with three kind of semantic segmentation source, i.e. the ground truth label (GT) and SOTA semantic segmentation method SEEM [16] with two backbones: the Large FocalNet [47] (SEEM-L) and the Tiny one (SEEM-T). SEEM-L is more accurate than SEEM-T in semantic segmentation. For detector-free matching, we choose the original ASpan [7], QuadT [11], LoFTR [10] and COTR [37]. For detector-based matching, we select SuperPoint [5] with nearest neighbor matching (SP+NN) and with the learning matching SuperGlue [30] (SP+SG). The recent SOTA ASLFeat+NN [9] and ALike+NN [8] are compared as well.
@@
2) Evaluation protocol. Following [23, 8], we report the Mean Matching Accuracy (MMA@i) in percentage under integer thresholds $i\in[1,3]$ of each method and the number of matches is set as 500 for each method.
@@
3) Results. The MMA values are reported in Tab. V, with improvement achieved by our methods are shown in percentage. It is evident that the SGAM significantly enhances the matching precision of all detector-free methods. It is worth noting that ScanNet serves as the training dataset for ASpan, QuadT, and LoFTR. Despite this, SGAM still demonstrates impressive accuracy improvement for these methods. On MatterPort3D dataset, SGAM exhibits substantial precision improvement and outperforms other methods by a noteworthy margin, thus underscoring the superiority of SGAM in challenging matching scenarios. Overall, our method consistently achieves significant improvements under various matching difficulties, highlighting the robustness and effectiveness of our approach. For each baseline, Tab. V also presents a comparison of different semantic inputs of our method. The ground truth labels are most precise, while SEEM, trained with COCO [48] labels, may introduce unidentified objects, resulting in a slight decrease in precision. SGAM using SEEM-L gets higher accuracy than using SEEM-T, demonstrating the matching accuracy increases with the semantic segmentation precision. Nevertheless, the semantic prior provided by the current semantic segmentation method is sufficient for our approach to achieve remarkable accuracy improvements in matching tasks. In sum, owing to the proper search space, which alleviates many matching challenges and provides high-resolution input for point matcher, SGAM notably improves the matching accuracy for SOTA detector-free matchers. We offer the visualization of qualitative results in Fig. 11.
Figure 4: Qualitative Comparison on ScanNet. The visual comparison between our method and two SOTA baselines. The wrong and correct matches under the same threshold are labeled respectively. The point matches we obtained possess much higher precision as well as uniform distribution.
@@
TABLE IV: Relative pose estimation Results (%) on KITTI360 dataset. We compare two different semantic inputs for our method: SGAM using ground truth (GT) and SGAM using SEEM-L.
Pose estimation AUC Sqe. 00 Seq. 03 Seq. 05
@@
ASpan[7] 61.19 77.64 87.95 68.01 83.00 91.20 57.38 75.57 87.16
@@
TABLE V: Relative pose estimation Results (%) on ScanNet-1500 benchmark. Our method obtains the semantic prior by SEEM-L. The best and second results are highlighted.
Pose estimation AUC ScanNet-1500 benchmark
@@
SP+SG+MKPC [14] 16.18-0.12% 34.11+0.92% 52.47+1.29%
DRC-Net[50] 7.70 17.90 30.50
MatchFormer[51] 24.30 43.90 61.40
PATS[15] 26.00 46.90 64.30
ASpan [7] 25.78 46.14 63.32
SEEM-L [16]+SGAM_ASpan 27.51+6.71% 48.01+4.05% 65.26+3.06%
@@
V-D Relative Pose Estimation
Accurate matches do not necessarily lead to accurate geometry, where point distribution is also important. Thus we next evaluate our method for relative pose estimation. The dataset used in our evaluation comprises both indoor and outdoor scenes. Specifically, we employ ScanNet and Matterport3D for indoor scenes. We sample $2\!\times\!1500$ image pairs from ScanNet (FD@5/10) and 500 image pairs from MatterPort3D to construct three difficulties. We also investigate the influence of different semantic inputs. Additionally, we compare our method with more SOTA methods [15, 32, 49, 51, 50, 14, 40] on the standard ScanNet-1500 benchmark [30], using SEEM-L to obtain the semantic segmentation results. Due to the static world assumption [1, 2] in our downstream tasks, we utilized three sequences in the KITTI360 dataset (Seq. 00, 03 and 05) with few moving objects, like pedestrians, for pose estimation estimation. In these sequences, we showcase the improvement achieved by our method across four baselines, using the semantic prior from both ground truth and SEEM-L results. For YFCC100M [46] dataset, we follow the previous work [30, 52] to construct pose estimation evaluation with 4k image pairs. We also use SEEM-L and SEEM-T to obtain semantic prior of this dataset for our method.

### v6

@@ III-D Framework Implementation
$\mathcal{P}_{g}^{C}=GMC_{PM}(\{\mathcal{A}_{i,\pi(i)},\mathcal{P}_{i}\}_{i}^{T},I_{0},I_{1},T_{SP}).$ (14)
Finally, the output point matches are merged by inside-area matches $\mathcal{P}_{i}\}_{i}^{T}$ and global matches $\mathcal{P}_{g}^{C}$, which possess both high matching accuracy and uniform spatial distribution. It is noteworthy the $PM$ we adopted can be any point matching method. Therefore, our SGAM is able to universally improve the accuracy of sparse, semi-dense and dense point matchers, as shown in our experiments.
IV Results
TABLE I: Value results (%) of MMA. We report MMA with three thresholds under various matching difficulties. Our SGAM is applied on four baselines. To show the impact of semantic accuracy to our method, we take three different semantic inputs: SGAM using ground truth (GT), SGAM using SEEM-L and SGAM using SEEM-T. The improvement achieved by SGAM is also reported in percentage, which is impressive to show the effectiveness of our methed.
Point Matching ScanNet: FD@5 ScanNet: FD@10 MatterPort3D
@@ IV Results
Sparse SP+SG [29] 37.54 63.06 76.15 24.40 43.32 57.57 13.77 21.66 29.95
GT+SGAM_SP+SG 41.74+11.18% 68.31+8.32% 81.46+6.98% 26.44+8.37% 44.96+3.79% 59.60+3.52% 15.94+15.80% 24.23+11.87% 32.86 +9.73%
SEEM-L [18]+SGAM_SP+SG 40.82+8.73% 66.68+5.74% 80.58+5.82% 25.65+5.14% 44.42+2.54% 59.42+3.21% 14.95+8.61% 23.36+7.85% 31.96+6.72%
SEEM-T [18]+SGAM_SP+SG 39.34+4.79% 65.31+3.56% 78.43+3.00% 25.31+3.74% 43.86+1.25% 58.65+1.87% 14.14+2.72% 22.86+5.55% 31.52+5.25%
Semi-Dense ASpan [10] 32.99 66.91 85.03 25.35 49.83 70.79 7.17 21.10 37.25
@@
SEEM-L+SGAM_ASpan 36.48+10.58% 70.70+5.66% 87.58+2.99% 27.15+7.09% 52.85+6.07% 73.76+4.19% 7.61+6.11% 23.98+13.66% 39.16+5.12%
SEEM-T+SGAM_ASpan 35.54+7.73% 69.44+3.79% 86.64+1.89% 26.81+5.78% 52.11+4.59% 72.84+2.89% 7.40+3.17% 22.51+6.69% 38.41+3.11%
QuadT [39] 32.79 70.40 88.31 22.67 56.92 78.46 7.44 23.97 41.72
@@
SEEM-L+SGAM_QuadT 37.02+12.91% 73.63+4.59% 89.30+1.12% 25.17+11.03% 60.55+6.38% 81.08+3.35% 7.91+6.41% 25.95+8.26% 43.05+3.17%
SEEM-T+SGAM_QuadT 36.35+10.85% 72.54+3.04% 88.40+0.10% 24.30+7.15% 59.38+4.32% 80.46+2.55% 7.86+5.63% 24.50+2.23% 42.88+2.76%
LoFTR [9] 30.49 65.33 83.51 17.85 46.78 67.90 9.50 22.08 36.07
@@
SEEM-L+SGAM_LoFTR 33.83+10.95% 70.05+7.23% 87.33+4.58% 18.85+5.60% 48.78+4.27% 68.90+1.47% 12.27+29.16% 27.20+23.22% 40.25+11.57%
SEEM-T+SGAM_LoFTR 33.17+10.55% 69.52+6.40% 86.71+3.84% 18.21+2.03% 47.45+1.42% 67.98+0.12% 11.47+20.77% 25.10+13.69% 38.45+6.59%
Dense COTR [35] 32.92 63.45 78.71 16.51 42.36 60.99 10.63 29.37 46.07
@@
SEEM-L+SGAM_COTR 36.54+11.00% 66.48+4.78% 81.04+2.96% 18.16+10.01% 44.54+5.15% 63.29+3.76% 11.73+10.40% 31.97+8.88% 48.70+5.71%
SEEM-T+SGAM_COTR 36.05+9.52% 65.89+3.85% 80.48+2.24% 17.60+6.60% 43.71+3.18% 62.50+2.47% 11.11+4.57% 31.30+6.58% 47.49+3.08%
IV-A Dataset
@@
To demonstrate the superiority of the A2PM framework and SGAM method, we first evaluate our methods on two different indoor datasets, ScanNet [42] and MatterPort3D [43]. Additionally, we investigated the robustness of our method in diverse semantic scenes by conducting experiments on the outdoor KITTI360 [44] and YFCC100M [45] dataset. First three datasets all offer ground truth semantic labels, which can be directly used as the semantic input of our method. Moreover, we also evaluated the practicability of our method by utilizing the input from recent semantic segmentation method, SEEM [18]. ScanNet contains numerous sequence images, and we selected image pairs with varying levels of difficulty based on the frame difference from its $scene\_0000$ to $scene\_0299$ to evaluate our method. We also compare with other SOTA methods on the standard ScanNet1500 benchmark [29], where semantic labels are achieved by SEEM. Due to the data collection settings of MatterPort3D, image pairs with overlap in this dataset have wide baseline and present challenging matching conditions. These conditions allow us to effectively showcase the performance of our method under difficult matching conditions. The KITTI360 dataset allows for the evaluation of driving scenes, which is widely-used for SLAM. The YFCC100M dataset contains internet images of architectural scenarios, which is widely-used for SfM.
@@
TABLE II: Relative pose estimation Results (%). The AUC of pose error on ScanNet (FD@5/10) and MatterPort3D with different thresholds are reported. Our SGAM is applied on four baselines. To show the impact of semantic accuracy to our method, we takes three different semantic inputs: SGAM using ground truth (GT), SGAM using SEEM-L and SGAM using SEEM-T. The improvement achieved by SGAM is also reported in percentage.
Pose Estimation ScanNet: FD@5 ScanNet: FD@10 MatterPort3D
@@
Sparse SP +SG[29] 67.46 76.46 86.61 53.11 64.47 73.58 16.39 29.54 37.61
GT+SGAM_SP+SG 69.20+2.58% 78.61+2.81% 88.72+2.44% 55.87+5.20% 66.87+3.72% 75.98+3.26% 17.93+9.40% 31.72+7.38% 39.31+4.52%
SEEM-L [18]+SGAM_SP+SG 68.61+1.70% 77.60+1.49% 87.44+0.96% 53.91+1.51% 65.46+1.54% 75.03+1.97% 17.15+4.64% 31.53+6.74% 38.51+2.39%
SEEM-T [18]+SGAM_SP+SG 68.33+1.30% 77.35+1.16% 87.13+0.60% 53.26+0.28% 65.11+0.99% 74.62+1.41% 16.85+2.81% 30.37+2.81% 37.95+0.90%
Semi-Dense ASpan [10] 70.73 77.41 80.19 58.51 70.42 79.84 18.35 27.81 43.98
@@
SEEM-L+SGAM_ASpan 72.32+2.24% 80.59+4.10% 85.20+6.24% 59.17+1.13% 73.02+3.69% 83.40+4.46% 19.74+7.56% 29.38+5.66% 45.52+3.51%
SEEM-T+SGAM_ASpan 71.84+1.56% 79.94+3.26% 83.32+3.90% 58.87+0.62% 72.02+2.27% 82.54+3.38% 18.64+1.56% 28.87+3.81% 44.34+0.83%
QuadT [39] 69.48 74.25 79.39 59.27 69.77 74.96 16.53 26.98 39.96
@@
SEEM-L+SGAM_QuadT 71.92+3.51% 75.60+1.81% 81.43+2.56% 61.78+4.23% 71.96+3.14% 76.52+2.08% 17.83+7.82% 27.25+1.00% 41.56+4.00%
SEEM-T+SGAM_QuadT 71.47+2.86% 75.06+1.09% 80.66+1.59% 60.97+2.87% 71.77+2.86% 75.80+1.13% 16.84+1.88% 27.10+0.44% 40.22+0.65%
LoFTR [9] 67.69 74.29 78.45 58.71 69.81 78.99 17.98 27.79 38.19
@@
SEEM-L+SGAM_LoFTR 70.25+3.79% 79.23+6.65% 83.93+6.99% 60.37+2.83% 71.02+1.74% 80.91+2.42% 18.04+0.31% 30.86+11.04% 42.13+10.32%
SEEM-T+SGAM_LoFTR 69.53+2.72% 78.03+5.03% 82.32+4.93% 59.01+0.50% 70.58+1.10% 79.78+1.01% 17.47-2.85% 30.00+7.94% 40.63+6.39%
Dense COTR [35] 66.91 74.11 78.48 51.92 63.36 72.55 17.80 25.08 34.08
@@
SEEM-L+SGAM_COTR 69.67+4.12% 78.15+5.46% 83.84+6.83% 52.80+1.70% 66.16+4.42% 78.78+8.59% 18.12+1.82% 26.99+7.62% 36.69+7.67%
SEEM-T+SGAM_COTR 69.40+3.72% 77.98+5.32% 83.07+5.85% 52.07+0.28% 65.91+4.03% 78.43+8.10% 17.93+0.71% 25.79+2.84% 35.79+5.01%
IV-B Point Matching
@@
We first conduct point matching experiments on ScanNet and MatterPort3D. For ScanNet, we construct two matching difficulties with image pairs under various Frame Differences (FD@5/10), each including 1500 image pairs. For the more challenging condition, 500 image pairs are sampled from the first 5 scenes in MatterPort3D.
@@
1) Compared methods. We compare the proposed SGAM method with various matching methods of all three types, including sparse method [7], semi-dense methods [10, 39, 9] and dense method [35]. Particularly, we combine SGAM with SOTA sparse method (SGAM_SP+SG [7, 29]), semi-dense methods (SGAM_ASpan [10], SGAM_QuadT [39] and SGAM_LoFTR [9]) and dense methods (SGAM_COTR [35]), to demonstrate the improvement we achieved. To evaluate the robustness against semantic input, we offers three semantic segmentation sources for SGAM, i.e. the ground truth label (GT) and semantic segmentation method SEEM [18] with two backbones: the Large FocalNet [46] (SEEM-L) and the Tiny one (SEEM-T). SEEM-L is more accurate than SEEM-T in semantic segmentation. The implementation details of our SGAM can be found in Sec. VI-A of the appendix.
@@
2) Evaluation protocol. Following [22, 6], we report the Mean Matching Accuracy (MMA@i) in percentage under integer thresholds $i\in[1,3]$ of each method. This metric indicates the proportion of correct matches among all matches. The number of matches is set as 500 for each method.
@@
3) Results. The MMA values are reported in Tab. I, with improvement achieved by our methods are shown in percentage. It is evident that the SGAM significantly enhances the matching precision of all three kinds of matching methods on both datasets, highlighting the robustness and effectiveness of our approach. Specifically, although ScanNet serves as the training dataset for ASpan, QuadT, and LoFTR, SGAM still demonstrates impressive accuracy improvement for these methods. On MatterPort3D dataset, SGAM also exhibits substantial precision improvement, thus underscoring the superiority of SGAM in challenging matching scenarios. For each baseline, Tab. I also presents a comparison of different semantic inputs of our method. The ground truth labels are most precise, while SEEM, trained with COCO [47] labels, may introduce unidentified objects, resulting in a slight decrease in precision. SGAM using SEEM-L gets higher accuracy than using SEEM-T, demonstrating the matching accuracy increases with the semantic segmentation precision. Nevertheless, the semantic prior provided by the current semantic segmentation method is sufficient for our approach to achieve remarkable accuracy improvements in matching tasks. In sum, owing to our improved search space, which alleviates many matching challenges and provides high-resolution input for point matchers, SGAM notably improves the matching accuracy. We offer the visualization in Fig. 8 of the appendix.
@@
TABLE III: Relative pose estimation Results (%) on KITTI360 dataset. We compare two different semantic inputs for our method: SGAM using ground truth (GT) and SGAM using SEEM-L.
Pose Estimation Sqe. 00 Seq. 03 Seq. 05
@@
Sparse SFD2 [16] 68.58 83.24 92.30 72.71 88.04 94.31 63.18 80.52 90.95
SP [7]+SG [29] 69.65 85.44 93.91 74.24 89.15 96.03 63.91 81.34 91.34
GT+SGAM_SP+SG 72.17+3.62% 86.31+1.02% 94.37+0.49% 75.36+1.51% 90.10+1.07% 97.26+1.28% 65.01+1.72% 82.78+1.77% 92.61+1.39%
SEEM-L[18]+SGAM_SP+SG 71.93+3.27% 86.06+0.73% 94.11+0.21% 75.23+1.33% 89.87+0.81% 96.99+1.00% 64.85+1.47% 82.53+1.46% 92.17+0.91%
Semi-Dense ASpan[10] 61.19 77.64 87.95 68.01 83.00 91.20 57.38 75.57 87.16
@@
TABLE IV: Relative pose estimation Results (%) on ScanNet1500 benchmark. Our method obtains the semantic prior by SEEM-L. The best and second results are highlighted.
Pose Estimation ScanNet1500 benchmark
@@
MKPC [13]+SP+SG 16.18-0.12% 34.11+0.92% 52.47+1.29%
SEEM-L [18]+SGAM_SP+SG 17.33+6.98% 34.77+2.87% 52.13+0.64%
Semi-Dense ASpan [10] 25.78 46.14 63.32
SEEM-L+SGAM_ASpan 27.51+6.71% 48.01+4.05% 65.26+3.06%
@@
IV-C Relative Pose Estimation
Accurate point matches do not necessarily lead to accurate geometry, as point distribution is also important. Thus we evaluate our method for relative pose estimation. The dataset used in our evaluation comprises both indoor and outdoor scenes. Specifically, we employ ScanNet and Matterport3D for indoor scenes. We sample $2\!\times\!1500$ image pairs from ScanNet (FD@5/10) and 500 image pairs from MatterPort3D to construct three difficulties. We also investigate the influence of different semantic inputs. Additionally, we compare our method with more SOTA methods [15, 31, 13] on the standard ScanNet1500 benchmark [29], using SEEM-L to obtain the semantic segmentation results. For outdoor scenes, we use the KITTI360 [44] and YFCC100M [45] dataset. Due to the static world assumption [1, 2] in downstream tasks, we utilized three sequences in the KITTI360 dataset (Seq. 00, 03 and 05) with few moving objects, e.g. pedestrians, for pose estimation estimation. In these sequences, we showcase the improvement achieved by our method across five baselines, using the semantic prior from both ground truth and SEEM-L results. For YFCC100M dataset, we follow the previous work [29] to construct pose estimation evaluation with 4k image pairs. We also use SEEM-L and SEEM-T to obtain semantic prior of this dataset for our method.